The moon has a mass of 7.32 Γ— 1022π‘˜π‘” and a radius of 1609.4 π‘˜π‘š. Calculate the value of "𝑔" at the surface of the moon.


1
Expert's answer
2021-10-03T13:25:08-0400
"g_M=(6.67\cdot10^{-11})\frac{7.32\cdot10^{22}}{(1609.4\cdot10^{3})^2}=1.88\frac{m}{s^2}"